Hello, can someone please help me with this problem:
Use the properties of exponents to rewrite y=4e^{-0.5t} in the form y=a(1+r)^t or y=a(1-r)^t . Round the value of r to the nearest thousandth. Then find the percent rate of change to the nearest tenth of percent.
I have tried but I really just don't understand how to do it.
Please help, Thanks.

Answer by ikleyn(53765)

.

y = 4*e^(-0.5t) = 4*(e^(-0.5))^t.


e^(-0.5) = 2.71828^(-0.5) = 0.6065 = 1 - 0.3935.


Therefore,


y = 4*e^(-0.5t) = 4*(e^(-0.5))^t = 4*(1-0.3935)^t


is that presentation (that formula), you are looking for.

Solved, explained, answered and completed.